Consider a monopolist whose total cost function is TC = 40 + 4Q + Q2 and whose marginal cost function is MC = 4 + 2Q. The demand function for the firm’s good is P = 160 - 0.5Q. The firm optimizes by producing the level of output that maximizes profit or minimizes loss. If the firm uses a uniform pricing strategy, then the price elasticity of demand (ED) at the profit maximizing price is equal to:
0.5
1.0
2.6
4.5
5.2
1
Expert's answer
2014-11-27T12:59:03-0500
TC = 40 + 4Q + Q2, MC = 4 + 2Q, P = 160 - 0.5Q.
If the firm uses a uniform pricing strategy, then the price elasticity of demand (ED) at the profit maximizing price is equal to the slope of the demand curve P = 160 - 0.5Q, so Ed = 0.5.
That's why the right answer is a) 0.5.